OpenStudy (anonymous):

Help with this problem? Please I have done majority of it. I just need help with the last part!

OpenStudy (anonymous):

Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 3:35 to 3:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 3π inches? Part 4: What is the coordinate point associated with this radian measure? You must show all of your work.

OpenStudy (anonymous):

I have done Part 1 and 2

OpenStudy (anonymous):

Part 1: minutes between 3:35 and 3:55 is 20 minutes/60 minutes = 1/3 * 2pi = 2.093 or 2Pi/3 radians. Part 2: there are 2 pi radians in a circle and the minute hand moves 20 minutes = 1/3 of the circle = 120 degrees = 120pi/180 radians = 2pi/3 radians formula length of arc = radius * angle in radians = 4 * 2pi/3 = 8pi/3

OpenStudy (anonymous):

I need help with Part 3 and 4 please

OpenStudy (amistre64):

what is the circumference of a unit circle?

OpenStudy (amistre64):

might not be reading it right we have a 4inch radial arm ... does the tip of it move along an arc of 3pi inches?

OpenStudy (amistre64):

 if so, then what portion of the circumference equates to 3pi? C = 2pi r kC = k 2pi r 3pi = k 2pi r solve for k

OpenStudy (amistre64):

we are a kth part of the original circumference so a kth part of a full circle is a radian measure if 2pi k

OpenStudy (anonymous):

k= √ 3r/r, -√3r/r

OpenStudy (anonymous):

Can anyone help with Part 4 please?

OpenStudy (amistre64):

3pi = k 2pi r 3 = 8k k = 3/8 so out of 2pi radians in a circle, we have moved 3/8 of 2pi assuming iver read the question correctly

OpenStudy (amistre64):